two jets are traveling towards each other and are 4000 km apart. The rate of one jet is 100 km/h faster than the rate of the other. if the jets pass each other after 2.5 hours, what is the rate of the faster jet?: two jets are traveling towards each other and are 4000 km apart. The rate of one jet is 100 km/h faster than the rate of the other. if the jets pass each other after 2.5 hours, what is the rate of the faster jet?
Answer by Alan3354(1449)

two jets are traveling towards each other and are 4000 km apart. The rate of one jet is 100 km/h faster than the rate of the other. if the jets pass each other after 2.5 hours, what is the rate of the faster jet?
-------------------------
The relative speed of one jet to the other is (2v + 100) km/hr, and is equal to the distance over the time, 4000/2.5 = 1600
2v+100 = 1600
2v = 1500
v = 750
The faster jet moves at 750+100 = 850 km/hour.